# Relaxation Fluctuations of Correlation Functions: Spin and Random Matrix Models

### v2

@@ VI Conclusion and Outlook
It is however to be noted that all the studies reported here are far from complete and general. All the results and the behaviour observed are with the specific choice of projection operators, which were originally chosen for their well-defined classical limits. A generalization to other operators, not necessarily projection operators is thus desirable. In these studies, we have restricted to mostly the case of $f=\frac{1}{2}$ and it would be desirable to study the variation of these quantities with partition $f$. As the quantity $c_{\text{avg}}$ has been argued to be a generalization of the IPR, it is thus of much interest to further explore its features. For the integrable cases, it is known that the coefficients are non-universal and show dependence on $f$, it would be important to study this dependence further. We also believe that the methods described here can be used to study in detail and quantify chaos in various other systems of present interest such as the Sachdev-Ye-Kitaev model ( see Rosenhaus (2019); Chowdhury et al. (2022) for a comprehensive review) and their sparse variants Xu et al. (2020b); Tezuka et al. (2023); Anegawa et al. (2023); Caceres et al. (2021); Cáceres et al. (2023).
Acknowledgements: The author would like to thank Masaki Tezuka for the critical reading of the draft and useful suggestions to shape the paper well. Numerical computations were performed using the computational facilities of the Yukawa Institute for Theoretical Physics. The Yukawa Research Fellowship of the author is supported by the Yukawa Memorial Foundation and JST CREST (Grant No. JPMJCR19T2).
@@ Appendix A Derivation for GUE
In this appendix, we provide the numerics that allow us to deduce the results for the GUE case. In Fig.(12) we provide the results for $c_{\text{avg}}$ and $\sigma_{\text{c}}^{2}$ for various values of $f$ with system sizes $D=128(2048),256(1024),512(512),1024(128),2048(64)$ ,4096(32),4096(32), 4096 ( 32 ), where the numbers in the bracket denote the number of samples. It is important to note that similar to the case of GOE we would expect our results to be universal in the sense that the results are independent of the partition $f$.
@@
If the point $z_{i}\in R^{D}$ are uniformly distributed on the surface of the sphere, then what is the resulting distribution?
@@
Using these properties, the distribution of Eq. (39) thus is $B(\frac{1}{2},\frac{n-1}{2})$. The required expectation value $\mathbb{E}{z_{1}^{4}}$ is then the second moment, $\mathbb{E}(Z_{1}^{2})=\mathbb{E}(z_{1}^{4})$, of $B(\frac{1}{2},\frac{n-1}{2})$. The PDF of the distribution is given as
@@
$\mathbb{E}(Z^{2})=\mathbb{E}(z^{2})=\frac{3}{D(D+2)}$ (42)
@@ Appendix B More results on Ising model
We now consider another set of parameters called the Kim-Huse parameters Kim and Huse (2013b). For the mixed field Ising Hamiltonian given by Eq (20)

### v3

@@ VI Conclusion and Outlook
It is however to be noted that all the studies reported here are far from complete and general. All the results and the behaviour observed are with the specific choice of projection operators, which were originally chosen for their well-defined classical limits. A generalization to other operators, not necessarily projection operators is thus desirable. In these studies, we have restricted to mostly the case of $f=\frac{1}{2}$ and it would be desirable to study the variation of these quantities with partition $f$. In the case of integrable cases, it is known that the coefficients in the scaling of $\sigma_{c}$ are non-universal and show dependence on $f$ Lakshminarayan (1997), it would be important to study this dependence further. As the quantity $c_{\text{avg}}$ has been argued to be a generalization of the IPR, it is thus of much interest to further explore its features in systems exhibiting localisation. We also believe that the methods described here can be used to study in detail and quantify chaos in various other systems of present interest such as the Sachdev-Ye-Kitaev model ( see Rosenhaus (2019); Chowdhury et al. (2022) for a comprehensive review) and their sparse variants Xu et al. (2020b); Tezuka et al. (2023); Anegawa et al. (2023); Caceres et al. (2021); Cáceres et al. (2023).
Acknowledgements: The author would like to thank Masaki Tezuka for the critical reading of the draft and useful suggestions to shape the paper well. The author would also like to thanks both the anonymous referees whose valuable comments improved the paper and made it more self contained. Numerical computations were performed using the computational facilities of the Yukawa Institute for Theoretical Physics. The Yukawa Research Fellowship of the author is supported by the Yukawa Memorial Foundation and JST CREST (Grant No. JPMJCR19T2).
@@ Appendix A Derivation for GUE
In this appendix, we provide some numerical and analytical results results for the GUE case. In Fig.(12) we provide the results for $c_{\text{avg}}$ and $\sigma_{\text{c}}^{2}$ for various values of $f$ with system sizes $D=128(2048),256(1024),512(512),1024(128),2048(64)$ ,4096(32),4096(32), 4096 ( 32 ), where the numbers in the bracket denote the number of samples used for averaging. It is important to note that similar to the case of GOE we would expect our results to be universal in the sense that the results are independent of the partition $f$.
@@
If the collection of points $\{z_{i}\}\in R^{D}$ are uniformly distributed on the surface of the sphere, then what is the resulting distribution?
@@
Using these properties, the distribution of Eq. (40) thus is Beta distribution $f_{Z}(\frac{1}{2},\frac{n-1}{2})$. The required expectation value $\mathbb{E}{(z_{1}^{4})}$ is then the second moment, $\mathbb{E}(Z_{1}^{2})=\mathbb{E}(z_{1}^{4})$, of $f_{Z}(\frac{1}{2},\frac{n-1}{2})$. The PDF of the distribution is given as
@@
$\mathbb{E}(Z^{2})\equiv\mathbb{E}(z^{4})=\frac{3}{D(D+2)}$ (43)
@@ Appendix B More results on Ising model
We now consider another set of parameters called the Kim-Huse parameters Kim and Huse (2013). For the mixed field Ising Hamiltonian given by Eq (21)